Find the differential equation of the family of curves x2+y2+2ax+2by+c=0, where a, b, c are parameters.


1
Expert's answer
2021-05-30T23:15:15-0400

"\displaystyle

x^2+y^2+2ax+2by+c=0\\

2x + 2yy' + 2a + 2by' = 0\\

y'(2b + 2y) = -(2a + 2x)\\

y' = -\frac{(a + x)}{(b + y)}"